# Bidimensional Exploration of the warm-Temperature Ionised gaS (BETIS)

### v1

@@ 3 The complex morphology of the eDIG
![Refer to caption](extracted/5688766/IC1553_eDIG_map_SII_NII.jpg)
![Refer to caption](extracted/5688766/IC1553_eDIG_map_OIII_OI.jpg)
@@
![Refer to caption](extracted/5688766/lines_z_IC1553.jpg)
3.1 Ionisation structure of the eDIG
@@ 4.1 The linkage between the galactic plane and the eDIG
Indeed, the connection between the distribution and the ionisation structure of the eDIG in the eBETIS galaxies (see Sec. 3.1) demonstrates a clear correlation between the morphology of halo emission and the distribution of SF regions and complexes in the underlying galaxy disks. Furthermore, the detection of extended eDIG in the halos of the eBETIS sample indicates that even low-mass galaxies with modest SFR (Table 2) are capable of producing a pervasive layer of extraplanar diffuse ionised emission. This finding contradicts previous assumptions that halos are a rare phenomenon among normal spirals (Hummel et al., 1991), a bias likely caused by the insufficient sensitivity of earlier observations.
@@ 4.2 The ionisation mechanism of the eDIG
Long-standing issues regarding the specifics of OB-star-driven ionisation remain prevalent in the literature. A notable characteristic of the DIG that proves difficult to explain through OB star photoionisation is its optical emission-line spectrum. Should the eDIG be ionised by radiation escaping from the midplane H II regions, the spectrum of the ionising radiation would be expected to resemble that of the OB star (or the SED dominated by such stars). Consequently, one would anticipate similar emission-line spectra in the eDIG as in the H II regions. Nonetheless, certain observed optical emission line ratios are not explicable by pure photoionisation from OB stars. The distinctive emission-line ratios observed in the eDIG, namely [N II] λ6584/H , [S II] λ6717/H , [O III] λ5007/H and [O I] λ6300/H , are elevated compared to those observed in H II regions (Rand, 1997, 1998; Collins & Rand, 2001; Haffner et al., 2009b; Seon, 2009), as illustrated in Fig. 3 and elaborated upon in Sec. 3.1 for the eBETIS sample, indicating the relative significance of additional heating processes.
@@
Relative temperatures can be explore by analysing the emission line ratios of the forbidden lines with respect to the H-recombination emission. The line intensities of [N II] and [S II] relative to H probe the ionisation fractions () and elemental abundances of ionised the gas, showing a strong dependence on the electron temperature (Te). These parameters relate with the line intensity ratios as follows:
@@
Haffner et al. (1999) and Madsen et al. (2006) used these relationships to build diagnostic diagrams to estimate both Te and S+/S from observations of [N II] /H and [S II] /H . Given that H and N have similar first ionisation potentials (13.56 eV and 14.5 eV, respectively) and a weak charge-exchange reaction, and assuming that N is not doubly ionised, then N+/N0 H+/H0 vary little within the DIG, so that variations in [N II] /H essentially traces variations in Te (Sembach et al., 2000; Otte et al., 2002; Haffner et al., 2009b). Figure 5 shows the distribution of [N II] λ 6583/H vs. [S II] λ 6717/H line ratios in function of —z— for IC1553. Circles are the integrated along the major axis for a certain z. Stars, triangles and squares are bins at MAD = -2.5, 0 and 3 kpc respectively. We use the equations 1 and 2 to show the predicted ratios for a constant Te (6, 7, 8 and 9 in units of 103 K) and S+/S (0.25, 0.5 and 0.8). We assume solar abundances: N/H = 6.8 and S/H = 1.3 (Asplund et al., 2021).
@@
Nevertheless, the behaviour of the line ratios shown in Fig. 3 (and related figures in the appendices) reveals that the ionisation of the eDIG is highly dependent on the distribution of H II regions in the disc (Sec. 3). The [N II] /H , [S II] /H , and [O III] /H ratios attain their highest values at the greatest distances from the midplane and in regions with lower . This observation is incompatible with a smooth distribution of HOLMES throughout the thick disc and lower halo. Under the HOLMES scenario, higher [O III] /H ratios would be expected in the vicinities of the thick disc, along with a smooth distribution of characteristic high ionisation line ratios within the halos. Furthermore, the few high [O III] /H ratio bins near the galactic planes of the galaxies are clearly associated with SF regions within the discs, as indicated by the emission-line maps of eBETIS (see Fig. 2 and related figures in the appendices).
@@
![Refer to caption](extracted/5688766/models_05.jpg)
![Refer to caption](extracted/5688766/models_03.jpg)
![Refer to caption](extracted/5688766/IC1553shocks_Z0.02_q1e7_0.6.jpg)
@@ 4.3 Hybrid SF and shocks ionisation mechanism of the eDIG
Shocks induced in the ISM by feedback mechanisms, such as supersonic winds originating from high-level SF regions, have been proposed as a significant source of heating for the eDIG333Note that tidal interactions and AGN jets can also produce shock ionisation in the halos of galaxies, with similar effects as discussed in the text (e.g., Dopita & Sutherland 1995; Simpson et al. 2007; Rich et al. 2011; Ho et al. 2016; Molina et al. 2018). (Chevalier & Clegg, 1985; Rand, 1998; Collins & Rand, 2001; Bland-Hawthorn et al., 2007). In this scenario, galactic winds are driven by the stellar winds and subsequent SNe of massive stars. These stellar winds create kpc-sized cavities in the ambient ISM on timescales of several million years (e.g. Mayya et al., 2023). Subsequently, these cavities are overrun by SN remnants (SNRs) from the most massive progenitor stars. Both processes occur at velocities exceeding the local sound speed, , in the ISM, thereby generating strong shocks.
@@
![Refer to caption](extracted/5688766/IC1553_SII_OI_GCD.jpg)
![Refer to caption](extracted/5688766/IC1553shocks_0.5.jpg)
The role of fast shocks as a potential alternative ionisation mechanism for the eDIG, as well as the impact of these filaments in the halo, has been studied by several authors (Rand, 1998; Collins & Rand, 2001; Levy et al., 2019; Dirks et al., 2023). For example, Martin 1997 studied the excitation of the DIG in 14 SF dwarf galaxies, arriving to the conclusion that photoionisation does dominate the excitation of the DIG, and that shocks are only being invoked as a secondary signal in gas with very low surface brightness, require the largest relative contribution from shocks, up to 30-50% of the total emission.
@@
For instance, Figure 7 shows the BPT diagram for all the eDIG bins for IC1553444Note that these results do not concur with the diagnosis of Rautio et al. (2022) for this galaxy, whose BPT diagrams are predominantly populated within the star formation regime. However, they align more closely with the subsequent study by Dirks et al. (2023). The locus in all three panels of the BPT diagram corresponds with the findings of this latter study, being proximate to the Kewley et al. (2001) line for [S II] /H , and achieving values up to 0 (in log scale) for [O I] /H . Nevertheless, in comparison to both studies, we observe a greater number of data points in the high excitation regimes for [S II] /H and [O I] /H , corresponding to bins at larger distances. We attribute this to our enhanced adaptive binning method, which enables sampling data with S/N ¿ 2 at greater distances, and to our exclusion of bins situated within the galactic plane, as discussed in section 3.. The models plotted represent the predicted fluxes for a combination of fast shocks with pre-ionisation (black curve lines) and with front shock only (blue curve lines) weighted with = 0.4, plus a photoionisation model correspondent to Z = Z⊙ and q = 107 cm/s, for instantaneous burst at 0 Myr and = 100 cm-3. Thus, the fluxes predicted for this particular hybrid models has the form: . Solid curves correspond to the shocks winds from 200 to 1000 km/s, and dashed curves represents magnetic field intensities from 0.0001 to 10 Gcm3/2. The parameters were selected to best fit the points in the BPT diagram, particularly the [S II] /H and [O I] /H diagnostics, as they are more sensitive to changes in ionisation regimes. The points in the [S II] /H and [O I] /H diagnostics tend to deviate more from photoionisation models than the [N II] /H diagnostic if an additional source of ionisation, such as AGN or fast shocks, is present (Dirks et al., 2023).
@@ 5.1 The radial variation of shocks contribution and the biconical structure of IC1553
![Refer to caption](extracted/5688766/IC1553shocks_Z0.02_q1e7_0.6_0.5.jpg)
![Refer to caption](extracted/5688766/IC1553shocks_Z0.02_q1e7_0.8_3.jpg)
Secondly, while the structure in the [N II] /H emission-line map appears to symmetrically originate from the disk at the galaxy’s minor axis (MAD kpc), the [S II] /H map reveals a less defined biconical structure with a discernible offset at MAD kpc. This same structure is observed in the [S II] /[N II] map, but with an opposite offset at MAD kpc. Additionally, the upper part of the [O I] /H map hints at this structure, displaying an offset at MAD kpc relative to [N II] /H . Most strikingly, the [O III] /H line-ratio map reveals no such structure on either side of the midplane. The high [O III] emission relative to H is a key observational feature of galactic outflows, regardless of whether its origin is due to SF (e.g. Tenorio-Tagle & Bodenheimer, 1988; Norman & Ikeuchi, 1989) or AGNs (e.g. López-Cobá et al., 2020; González-Díaz et al., 2023). The [O III] /H map of IC1553 exhibits the typical increase of the ratio with distance from the midplane in a homogeneous manner, independent of radial distance, but without any characteristic signature of an outflow at MAD . Therefore, based on the information provided by the 2D distribution of the eDIG, there is no morphological congruence among the different emission-line maps to definitively confirm the existence of an outflow in IC1553.
@@
galaxy | DIG morphology | MAD, z | LeDIG(H ) | log SFRDisc | ||
(kpc) | (erg/s) | (M⊙/yr) | ||||
@@ 6 Summary and conclusions
To account for the [O III] /H and [O I] /H ratios, various ionisation mechanisms have been considered. One such mechanism, recently highlighted in the literature, involves photoionisation by a hard spectrum emanating from the high-temperature end of the white dwarf distribution and the central stars of planetary nebulae; the low-mass, hot, and evolved (post-AGB) stars known as HOLMES. Nevertheless, the correlation between the ionisation structure of the eDIG and the H II regions distribution observed in this study does not support a homogeneous distribution of HOLMES across the thick disc and lower halo. This inconsistency, combined with the unreliability of EWHαas a proxy commonly used in the literature to differentiate star-forming regions from HOLMES, effectively negates HOLMES as a plausible ionisation mechanism for the eDIG.
@@
A prime example of this phenomenon is IC1553. In the [N II] /H emission line map, it reveals a biconical central structure within the halo, previously interpreted as indicative of outflows or superbubbles. Nonetheless, at the bicone’s location (MAD = 0), the galaxy’s disk is expected to exhibit a lower specific star formation rate (sSFR) along the line-of-sight due to the presence of the galaxy’s bulge, in contrast to the intense emission from the bright adjacent SF regions at the edge of the disk, as observed in projection, exhibiting typical H II region (low) emission values. This scenario is more pronounced in the [S II] /H and [O I] /H maps, where the background emission conspicuously permeates the halo with typical extended eDIG values. These observations, coupled with the absence of any structure in H , [O III] /H , the consistent [S II] /[N II] ratio, and the lack of spatial correlation with potential ionisation sources within the disk, indicate that this bicone is an optical artefact resulting from the projection effect relative to the line-of-sight of the high-level SF regions in the galaxy’s disk, in relation to a pervasive background eDIG emission in the halo.
@@ Acknowledgements.
- González-Díaz et al. (2023) González-Díaz R., et al., 2023, arXiv e-prints, p. arXiv:2311.14254
@@ Appendix A PGC28308
![Refer to caption](extracted/5688766/PGC28308_eDIG_map_SII_NII.jpg)
![Refer to caption](extracted/5688766/PGC28308_eDIG_map_OIII_OI.jpg)
@@
![Refer to caption](extracted/5688766/lines_z_PGC28308.jpg)
![Refer to caption](extracted/5688766/lines_GCD_up_PGC28308.jpg)
![Refer to caption](extracted/5688766/PGC28308shocks_Z0.02_q1e7_0.5.jpg)
![Refer to caption](extracted/5688766/PGC28308_S_T_GCD.jpg)
![Refer to caption](extracted/5688766/PGC28308_SII_OI_GCD.jpg)
@@ Appendix B PGC30591
![Refer to caption](extracted/5688766/PGC30591_eDIG_map_SII_NII.jpg)
![Refer to caption](extracted/5688766/PGC30591_eDIG_map_OIII_OI.jpg)
@@
![Refer to caption](extracted/5688766/PGC30591shocks_Z0.02_q1e7_0.6.jpg)
![Refer to caption](extracted/5688766/PGC30591_S_T_GCD.jpg)
![Refer to caption](extracted/5688766/PGC30591_SII_OI_GCD.jpg)
@@ Appendix C ESO544-27
![Refer to caption](extracted/5688766/ESO544-27_eDIG_map_SII_NII.jpg)
![Refer to caption](extracted/5688766/ESO544-27_eDIG_map_OIII_OI.jpg)
@@
![Refer to caption](extracted/5688766/lines_z_ESO544-27.jpg)
![Refer to caption](extracted/5688766/lines_GCD_up_ESO544-27.jpg)
@@
![Refer to caption](extracted/5688766/ESO544-27shocks_Z0.02_q1e7_0.5.jpg)
![Refer to caption](extracted/5688766/ESO544-27_S_T_GCD.jpg)
![Refer to caption](extracted/5688766/ESO544-27_SII_OI_GCD.jpg)

### v2

@@ 4.1 The linkage between the galactic plane and the eDIG
Indeed, the connection between the distribution and the ionisation structure of the eDIG in the eBETIS galaxies (see Sec. 3) demonstrates a clear correlation between the morphology of halo emission and the distribution of SF regions and complexes in the underlying galaxy disks. Furthermore, the detection of extended eDIG in the halos of the eBETIS sample indicates that even low-mass galaxies with modest SFR (Tables 1 and 6) are capable of producing a pervasive layer of extraplanar diffuse ionised emission. This finding contradicts previous assumptions that halos are a rare phenomenon among normal spirals (Hummel et al., 1991), a bias likely caused by the insufficient sensitivity of earlier observations.
@@ 4.2 The ionisation mechanism of the eDIG
Long-standing issues regarding the specifics of OB-star-driven ionisation remain prevalent in the literature. A notable characteristic of the DIG that proves difficult to explain through OB star photoionisation is its optical emission-line spectrum. Should the eDIG be ionised by radiation escaping from the midplane H II regions, the spectrum of the ionising radiation would be expected to resemble that of the OB star (or the SED dominated by such stars). Consequently, one would anticipate similar emission-line spectra in the eDIG as in the H II regions. Nonetheless, certain observed optical emission line ratios are not explicable by pure photoionisation from OB stars. The distinctive emission-line ratios observed in the eDIG, namely [N II] λ6584/H , [S II] λ6717/H , [O III] λ5007/H and [O I] λ6300/H , are elevated compared to those observed in H II regions (Rand, 1997, 1998; Collins & Rand, 2001; Haffner et al., 2009b; Seon, 2009), as illustrated in Fig. 3 and elaborated upon in Sec. 3 for the eBETIS sample, indicating the relative significance of additional heating processes.
@@
Relative temperatures can be explore by analysing the emission line ratios of the forbidden lines with respect to the H-recombination emission. The line intensities of [N II] and [S II] relative to H probe the ionisation fractions () and elemental abundances of ionised the gas, showing a strong dependence on the electron temperature (Te). These parameters relate with the line intensity ratios as follows (Haffner et al., 1999; Osterbrock & Ferland, 2006):
@@
We assume that usually the 100% of the hydrogen and the 80% of the nitrogen are ionised in the DIG, varying little within the DIG and basically tracing the Te (Sembach et al., 2000; Otte et al., 2002; Haffner et al., 2009b). We also assume solar abundances: N/H = 6.8 and S/H = 1.3 (Asplund et al., 2021).
Haffner et al. (1999) and Madsen et al. (2006) used these relationships to build diagnostic diagrams to estimate both Te and S+/S from observations of [N II] /H and [S II] /H . Given that H and N have similar first ionisation potentials (13.56 eV and 14.5 eV, respectively) and a weak charge-exchange reaction, and assuming that N is not doubly ionised, then N+/N0 H+/H0 vary little within the DIG, so that variations in [N II] /H essentially traces variations in Te (Sembach et al., 2000; Otte et al., 2002; Haffner et al., 2009b). Figure 5 shows the distribution of [N II] λ 6583/H vs. [S II] λ 6717/H line ratios in function of —z— for IC1553. Circles are the integrated along the major axis for a certain z. Stars, triangles and squares are bins at MAD = -2.5, 0 and 3 kpc respectively. We use the equations 1 and 2 to show the predicted ratios for a constant Te (6, 7, 8 and 9 in units of 103 K) and S+/S (0.25, 0.5 and 0.8).
@@
The increase in the [S II] /H and [N II] /H ratios with distance from the midplane is commonly used to support HOLMES as an ionisation source in the eDIG (Flores-Fajardo et al., 2011; Rautio et al., 2022), as well as the radial behaviour of these lines in low-inclination spiral galaxies (Belfiore et al., 2022). Nevertheless, the behaviour of the line ratios shown in Fig. 3 (and related figures in the appendices) reveals that the ionisation of the eDIG is highly dependent on the distribution of H II regions in the disc (Sec. 3). The [N II] /H , [S II] /H , and [O III] /H ratios attain their highest values at the greatest distances from the midplane and in regions with lower . This observation is incompatible with a smooth distribution of HOLMES throughout the thick disc and lower halo. Under the HOLMES scenario, higher [O III] /H ratios would be expected in the vicinities of the thick disc, along with a smooth distribution of characteristic high ionisation line ratios within the halos. Furthermore, the few high [O III] /H ratio bins near the galactic planes of the galaxies are clearly associated with SF regions within the discs, as indicated by the emission-line maps of eBETIS (see Fig. 2 and related figures in the appendices). Then, the behaviour of the [N II] /H , [S II] /H , and [O III] /H ratios alone is not sufficient evidence to conclusively support HOLMES as an ionisation source.
@@ 4.3 Hybrid SF and shocks ionisation mechanism of the eDIG
Shocks induced in the ISM by feedback mechanisms, such as supersonic winds originating from high-level SF regions, have been proposed as a significant source of heating for the eDIG444Note that tidal interactions and AGN jets can also produce shock ionisation in the halos of galaxies, with similar effects as discussed in the text (e.g., Dopita & Sutherland 1995; Simpson et al. 2007; Rich et al. 2011; Ho et al. 2016; Molina et al. 2018). (Chevalier & Clegg, 1985; Rand, 1998; Collins & Rand, 2001; Bland-Hawthorn et al., 2007). In this scenario, galactic winds are driven by the stellar winds and subsequent SNe of massive stars. These stellar winds create kpc-sized cavities in the ambient ISM on timescales of several million years (e.g. Mayya et al., 2023). Subsequently, these cavities are overrun by SN remnants (SNRs) from the most massive progenitor stars. Both processes occur at velocities exceeding the local sound speed, , in the ISM, thereby generating strong shocks.
@@
The role of fast shocks as a potential alternative ionisation mechanism for the eDIG, as well as the impact of these filaments in the halo, has been studied by several authors (Rand, 1998; Collins & Rand, 2001; Levy et al., 2019; Dirks et al., 2023). For instance, Martin (1997) studied the excitation of the DIG in 14 star-forming dwarf galaxies and concluded that photoionisation due to star formation is the dominant mechanism. Shocks were only considered as a secondary ionisation source for gas with very low surface brightness. In this scenario, shocks could contribute up to 30-50% of the total ionisation budget.
@@
For instance, Figure 7 shows the BPT diagram for all the eDIG bins for IC1553555Note that these results do not concur with the diagnosis of Rautio et al. (2022) for this galaxy, whose BPT diagrams are predominantly populated within the star formation regime. However, they align more closely with the subsequent study by Dirks et al. (2023). The locus in all three panels of the BPT diagram corresponds with the findings of this latter study, being proximate to the Kewley et al. (2001) line for [S II] /H , and achieving values up to 0 (in log scale) for [O I] /H . Nevertheless, in comparison to both studies, we observe a greater number of data points in the high excitation regimes for [S II] /H and [O I] /H , corresponding to bins at larger distances. We attribute this to our enhanced adaptive binning method, which enables sampling data with S/N ¿ 2 at greater distances, and to our exclusion of bins situated within the galactic plane, as discussed in section 3.. The models plotted represent the predicted fluxes for a combination of fast shocks with pre-ionisation (black curve lines) and with front shock only (blue curve lines) weighted with = 0.4, plus a photoionisation model correspondent to Z = Z⊙ and q = 107 cm/s, for instantaneous burst at 0 Myr and = 100 cm-3. Thus, the fluxes predicted for this particular hybrid models has the form: . Solid curves correspond to the shocks winds from 200 to 1000 km/s, and dashed curves represents magnetic field intensities from 0.0001 to 10 Gcm3/2. The parameters were selected to best fit the points in the BPT diagram, particularly the [S II] /H and [O I] /H diagnostics, as they are more sensitive to changes in ionisation regimes. The points in the [S II] /H and [O I] /H diagnostics tend to deviate more from photoionisation models than the [N II] /H diagnostic if an additional source of ionisation, such as AGN or fast shocks, is present (Dirks et al., 2023). The for fitting each BPT was determined by performing several trials with different , selecting the one that best fits the data points in the [S II] /H and [O I] /H diagnostics.
@@ 5.1 The radial variation of shocks contribution and the biconical structure of IC1553
Secondly, while the structure in the [N II] /H emission-line map appears to symmetrically originate from the disk at the galaxy’s minor axis (MAD kpc), the [S II] /H map reveals a less defined biconical structure with a discernible offset at MAD kpc. This same structure is observed in the [S II] /[N II] map, but with an opposite offset at MAD kpc. Additionally, the upper part of the [O I] /H map hints at this structure, displaying an offset at MAD kpc relative to [N II] /H . Most strikingly, the [O III] /H line-ratio map reveals no such structure on either side of the midplane. The high [O III] emission relative to H is a key observational feature of galactic outflows, regardless of whether its origin is due to SF (e.g. Tenorio-Tagle & Bodenheimer, 1988; Norman & Ikeuchi, 1989) or AGNs (e.g. López-Cobá et al., 2020; González-Díaz et al., 2024). The [O III] /H map of IC1553 exhibits the typical increase of the ratio with distance from the midplane in a homogeneous manner, independent of radial distance, but without any characteristic signature of an outflow at MAD . Therefore, based on the information provided by the 2D distribution of the eDIG, there is no morphological congruence among the different emission-line maps to definitively confirm the existence of an outflow in IC1553.
@@
| galaxy | DIG morphology | MAD, z | ||
@@ 6 Summary and conclusions
To account for the [O III] /H and [O I] /H ratios, various ionisation mechanisms have been considered. One such mechanism, recently highlighted in the literature, involves photoionisation by a hard spectrum emanating from the high-temperature end of the white dwarf distribution and the central stars of planetary nebulae; the low-mass, hot, and evolved (post-AGB) stars known as HOLMES. Nevertheless, the correlation between the ionisation structure of the eDIG and the H II regions distribution observed in this study does not support a homogeneous distribution of HOLMES across the thick disc and lower halo. This inconsistency, combined with the unreliability of EWHαas a proxy commonly used in the literature to differentiate star-forming regions from HOLMES, effectively negates HOLMES as a plausible ionisation mechanism for the eDIG. Consequently, the behaviour of the [N II] /H , [S II] /H , and [O III] /H ratios alone is not sufficient evidence to conclusively support HOLMES as an ionisation source. These behaviours should be carefully used in the analysis of the ionisation mechanisms of the (e)DIG.
@@
A prime example of this phenomenon is IC1553. In the [N II] /H emission line map, it reveals a biconical central structure within the halo, previously interpreted as indicative of outflows or superbubbles. Nonetheless, at the bicone’s location (MAD = 0), the galaxy’s disk is expected to exhibit a lower specific star formation rate (sSFR) along the line-of-sight due to the presence of the galaxy’s bulge. This contrasts with the intense emission from the bright adjacent star-forming regions at the edge of the disk, observed in projection, which exhibit typical H II region (low) emission values. This scenario is more pronounced in the [S II] /H and [O I] /H maps, where the background emission conspicuously permeates the halo with typical extended eDIG values. These observations, coupled with the absence of any structure in H , [O III] /H , the consistent [S II] /[N II] ratio, and the lack of spatial correlation with potential ionisation sources within the disk, indicate that this bicone is an optical artefact resulting from the projection effect relative to the line-of-sight of the high-level SF regions in the galaxy’s disk, in relation to a pervasive background eDIG emission in the halo.
@@ Acknowledgements.
- González-Díaz et al. (2024) González-Díaz R., et al., 2024, A&A, 687, A20